OpenStudy (anonymous):

is cosx=sqrt3/4 an identity

OpenStudy (anonymous):

That's what you mean?\[\cos(x)=\sqrt{\frac{ 3 }{4 }}\]

OpenStudy (anonymous):

answer please YES or NO?

OpenStudy (anonymous):

its not considered an "identity" but it is an important angle you should know.

OpenStudy (anonymous):